Write v as a linear combination of u1, u2, u3, where

(a) v = (4, −9, 2), u1 = (1, 2, −1), u2 = (1, 4, 2), u3 = (1, −3, 2);

(b) v = (1, 3, 2), u1 = (1, 2, 1), u2 = (2, 6, 5), u3 = (1, 7, 8);

(c) v = (1, 4, 6), u1 = (1, 1, 2), u2 = (2, 3, 5), u3 = (3, 5, 8);


1
Expert's answer
2022-01-25T12:18:35-0500

(a) Writing "~v" as a linear combination of "u_1, u_2," and "u_3" means finding "c_1," "c_2," and "c_3" such that "v=c_1u_1+c_2u_2+c_3u_3." This is equivalent to solving a system of linear equations


"\begin{bmatrix}
 1 & 1 & 1 \\
 2 & 4 & -3 \\
 -1 & 2 & 2
\end{bmatrix}\begin{bmatrix}
 c_1 \\
 c_2 \\
 c_3
\end{bmatrix}=\begin{bmatrix}
 4 \\
 -9 \\
 2
\end{bmatrix}"

Augmented matrix


"A=\begin{bmatrix}
 1 & 1 & 1 &\ \ 4\\
 2 & 4 & -3 & \ \ -9 \\
 -1 & 2 & 2 & \ \ 2
\end{bmatrix}"

"R_2=R_2-2R_1"


"\begin{bmatrix}
 1 & 1 & 1 &\ \ 4\\
 0 & 2 & -5 & \ \ -17 \\
 -1 & 2 & 2 & \ \ 2
\end{bmatrix}"

"R_3=R_3+R_1"


"\begin{bmatrix}
 1 & 1 & 1 &\ \ 4\\
 0 & 2 & -5 & \ \ -17 \\
 0 & 3 & 3 & \ \ 6
\end{bmatrix}"

"R_2=R_2\/2"


"\begin{bmatrix}
 1 & 1 & 1 &\ \ 4\\
 0 & 1 & -5/2 & \ \ -17/2 \\
 0 & 3 & 3 & \ \ 6
\end{bmatrix}"

"R_1=R_1-R_2"


"\begin{bmatrix}
 1 & 0 & 7/2 &\ \ 25/2\\
 0 & 1 & -5/2 & \ \ -17/2 \\
 0 & 3 & 3 & \ \ 6
\end{bmatrix}"

"R_3=R_3-3R_2"


"\begin{bmatrix}
 1 & 0 & 7/2 &\ \ 25/2\\
 0 & 1 & -5/2 & \ \ -17/2 \\
 0 & 0 & 21/2 & \ \ 63/2
\end{bmatrix}"

"R_3=2R_3\/21"


"\begin{bmatrix}
 1 & 0 & 7/2 &\ \ 25/2\\
 0 & 1 & -5/2 & \ \ -17/2 \\
 0 & 0 & 1 & \ \ 3
\end{bmatrix}"

"R_1=R_1-7R_3\/2"


"\begin{bmatrix}
 1 & 0 & 0 &\ \ 2\\
 0 & 1 & -5/2 & \ \ -17/2 \\
 0 & 0 & 1 & \ \ 3
\end{bmatrix}"

"R_2=R_2+5R_3\/2"


"\begin{bmatrix}
 1 & 0 & 0 &\ \ 2\\
 0 & 1 & 0 & \ \ -1 \\
 0 & 0 & 1 & \ \ 3
\end{bmatrix}"

So "c_1 = 2, c_2 = -1, c_3 = 3," and we can check that we can write


"v=2u_1-u_2+3u_3"

"=2\begin{bmatrix}
 1 \\
 2 \\
 -1
\end{bmatrix}-\begin{bmatrix}
 1 \\
 4 \\
 2
\end{bmatrix}+3\begin{bmatrix}
 1 \\
 -3 \\
 2
\end{bmatrix}=\begin{bmatrix}
 4 \\
 -9 \\
 2
\end{bmatrix}"

(b) Writing "~v" as a linear combination of "u_1, u_2," and "u_3" means finding "c_1," "c_2," and "c_3" such that "v=c_1u_1+c_2u_2+c_3u_3." This is equivalent to solving a system of linear equations


"\begin{bmatrix}
 1 & 2 & 1 \\
 2 & 6& 7 \\
 1 & 5 & 8
\end{bmatrix}\begin{bmatrix}
 c_1 \\
 c_2 \\
 c_3
\end{bmatrix}=\begin{bmatrix}
 1 \\
 3 \\
 2
\end{bmatrix}"

Augmented matrix


"A=\begin{bmatrix}
 1 & 2 & 1 &\ \ 1\\
 2 & 6 & 7 & \ \ 3 \\
 1 & 5 & 8 & \ \ 2
\end{bmatrix}"

"R_2=R_2-2R_1"


"\begin{bmatrix}
 1 & 2 & 1 &\ \ 1\\
 0 & 2 & 5 & \ \ 1 \\
 1 & 5 & 8 & \ \ 2
\end{bmatrix}"

"R_3=R_3-R_1"


"\begin{bmatrix}
 1 & 2 & 1 &\ \ 1\\
 0 & 2 & 5 & \ \ 1 \\
 0 & 3 & 7 & \ \ 1
\end{bmatrix}"

"R_2=R_2\/2"


"\begin{bmatrix}
 1 & 2 & 1 &\ \ 1\\
 0 & 1 & 5/2 & \ \ 1/2 \\
 0 & 3 & 7 & \ \ 1
\end{bmatrix}"

"R_1=R_1-2R_2"


"\begin{bmatrix}
 1 & 0 & -4 &\ \ 0\\
 0 & 1 & 5/2 & \ \ 1/2 \\
 0 & 3 & 7 & \ \ 1
\end{bmatrix}"

"R_3=R_3-3R_2"


"\begin{bmatrix}
 1 & 0 & -4 &\ \ 0\\
 0 & 1 & 5/2 & \ \ 1/2 \\
 0 & 0 & -1/2 & \ \ -1/2
\end{bmatrix}"

"R_3=-2R_3"


"\begin{bmatrix}
 1 & 0 & -4 &\ \ 0\\
 0 & 1 & 5/2 & \ \ 1/2 \\
 0 & 0 & 1 & \ \ 1
\end{bmatrix}"

"R_1=R_1+4R_3"


"\begin{bmatrix}
 1 & 0 & 0 &\ \ 4\\
 0 & 1 & 5/2 & \ \ 1/2 \\
 0 & 0 & 1 & \ \ 1
\end{bmatrix}"

"R_2=R_2-5R_3\/2"


"\begin{bmatrix}
 1 & 0 & 0 &\ \ 4\\
 0 & 1 & 0& \ \ -2 \\
 0 & 0 & 1 & \ \ 1
\end{bmatrix}"

So "c_1 = 4, c_2 = -2, c_3 = 1," and we can check that we can write


"v=4u_1-2u_2+u_3"



"=4\begin{bmatrix}
 1 \\
 2 \\
 1
\end{bmatrix}-2\begin{bmatrix}
 2 \\
 6\\
 5
\end{bmatrix}+\begin{bmatrix}
 1 \\
 7 \\
 8
\end{bmatrix}=\begin{bmatrix}
 1 \\
 3 \\
 2
\end{bmatrix}"

(c) Writing "~v" as a linear combination of "u_1, u_2," and "u_3" means finding "c_1," "c_2," and "c_3" such that "v=c_1u_1+c_2u_2+c_3u_3." This is equivalent to solving a system of linear equations


"\begin{bmatrix}
 1 & 2 & 3 \\
 1 & 3 & 5 \\
 2 & 5 & 8
\end{bmatrix}\begin{bmatrix}
 c_1 \\
 c_2 \\
 c_3
\end{bmatrix}=\begin{bmatrix}
 1 \\
 4 \\
 6
\end{bmatrix}"

Augmented matrix


"A=\begin{bmatrix}
 1 & 2 & 3 &\ \ 1\\
 1 & 3 & 5 & \ \ 4 \\
 2 & 5 & 8 & \ \ 6
\end{bmatrix}"

"R_2=R_2-R_1"

"\begin{bmatrix}
 1 & 2 & 3 &\ \ 1\\
 0 & 1 & 2 & \ \ 3 \\
 2 & 5 & 8 & \ \ 6
\end{bmatrix}"

"R_3=R_3-2R_1"

"\begin{bmatrix}
 1 & 2 & 3 &\ \ 1\\
 0 & 1 & 2 & \ \ 3 \\
 0 & 1 & 2 & \ \ 4
\end{bmatrix}"

"R_1=R_1-2R_2"

"\begin{bmatrix}
 1 & 0 & -1 &\ \ -5\\
 0 & 1 & 2 & \ \ 3 \\
 0 & 1 & 2 & \ \ 4
\end{bmatrix}"

"R_3=R_3-R_2"

"\begin{bmatrix}
 1 & 0 & -1 &\ \ -5\\
 0 & 1 & 2 & \ \ 3 \\
 0 & 0 & 0 & \ \ 1
\end{bmatrix}"

"R_1=R_1+5R_3"

"\begin{bmatrix}
 1 & 0 & -1 &\ \ 0\\
 0 & 1 & 2 & \ \ 3 \\
 0 & 0 & 0 & \ \ 1
\end{bmatrix}"

Then the system has no solution.

Therefore we cannot write "v" as a linear combination of "u_1, u_2, u_3."